# bar height=.6

### v1

@@ 4. Operational Carbon Mitigation
Intelligent Decisions. We propose a cognitive stack (kahneman11, ), which permits the separation of concerns and clean abstractions. The stack organizes power management into a fast, low-level reactive layer that is vertically coupled to a strategic, high-level deliberative layer. An agent monitors local job performance and resource utilization, optimizing its power requests to achieve its performance goals while accounting for global datacenter conditions and competition from other agents. For example, the reactive policy adjusts an individual processor’s power mode in response to fine-grained program phases while the deliberative policy ensures each processor’s adjustment anticipates other processors’ policies and the datacenter’s broader sustainability, safety, and stability goals.
@@
The cognitive stack leverages multi-agent game theory for dynamic decision making. Dynamism is important because computation exhibits time-varying behavior, and allocation decisions in the present should account for the past and anticipate the future. For example, consider a repeated game in which agents spend tokens for power. Each agent learns a policy for spending tokens, requesting power, and mapping power to datacenter resources. When carbon-free energy is scarce, the datacenter could offer more tokens to users that defer jobs or require more tokens from those who compute. How should agents spend tokens to maximize long-term performance when allocations in one time period affect those in an uncertain future? How should the datacenter price power to achieve sustainability or DR goals?
Risks. Some might wonder whether DR will be necessary given how datacenter operators are investing in renewable energy. Today, net zero claims rely on wind/solar energy investments that offset datacenter energy use, but carbon-free energy is too often generated at times and places that do not align with when and where computation happens such that ”net-zero” datacenters consume carbon-intensive energy in many hours of the year.
@@
Others might wonder whether DR already exists. Today’s narrow solutions focus on stability, requesting reduced use to avoid rare power emergencies, but even these suffer from incentive problems (wierman14, ). Researchers have studied how datacenters could participate in DR using simplifying assumptions (e.g., 20% of power is consumed by batch jobs, deferrable within a 24-hour window without loss of utility) (wierman14, ; goiri13, ; radovanovic21, ), but these solutions have limited scope.
@@ 5. Energy Economics
Economics and policy shape the solution space for carbon-efficient computing. Governments might implement carbon trading or incentives for low-carbon energy. The private sector might implement offset programs, leading to renewable energy purchase agreements and credits. Future demand response frameworks will require sophisticated markets that price electricity at its true social marginal cost and incentivize users to schedule computation accordingly. Although there is extensive literature on low-carbon policies for other industries (abrell19, ), economic analysis of policies specifically aimed at computing is relatively unexplored. In the near term, industry will benefit from policy-induced incentives when investing in renewable energy supply (e.g., renewable energy certificates). But as supply grows, industry must be able to monetize its flexibility in energy demand. Datacenters are often the largest consumers on the grid and we will study how their locally optimized decisions for net zero operations will affect other consumers and impact society.
Economics and policy shape the solution space for carbon-efficient computing. Governments might implement policy for carbon trading or energy incentives. The private sector might implement offset programs, leading to renewable energy purchase agreements and credits. Future demand response frameworks rely on incentives, which require sophisticated markets. Although there is extensive literature on low-carbon policies for other industries (abrell19, ), economic analysis for computing is relatively unexplored. In the near term, industry will receive credit when investing in renewable energy supply. But as supply grows, industry must respond to and receive credit for flexibility in their energy demand. Datacenters are often the largest consumers on the grid and we will study how their locally optimized decisions for net zero operations will affect other consumers and impact society. Furthermore, we will study how improved sustainability impact demand for computing.
@@
Given an unpriced environmental externality (hardin68, ), such as carbon, one might ask whether society is computing too much. What is the optimal amount of computing for society? Will more efficient algorithms and systems lead to so much demand for new computing applications that overall carbon will increase (jevons66, )? The Jevons Paradox states increased efficiency may not reduce demand for energy in the long run (and may even increase it). Prior research suggests, as a technology becomes more efficient, its use increases and produces rebound effects that range from 10% to 40%, reducing but not eliminating energy savings (gillingham16, ). But there has been no study of these effects for computing.
@@
We will estimate three types of rebound effects as technological efficiency lowers operating costs. First, direct effects arise when lower costs increase use of the technology. Datacenters likely exhibit strong direct effects as operators provision hardware to maximally use provisioned power (fan07, ); more efficient processors lead to datacenters with more processors. Second, indirect effects arise when lower costs increase use of other technologies. Quantifying indirect effects requires understanding substitutability and complementary between hardware components, which in turn depend on hardware capacity translates into software performance; more efficient processors may lead to datacenters with more memory as well. Finally, macroeconomic effects arise when lower costs encourage technology use for new applications. Efficient processors may scale the use of large language models for everyday tasks (e.g., search queries) rather than niche tasks (e.g., playing games).

### v2

@@ 4. Operational Carbon
Intelligent Decisions. A cognitive stack would permit the separation of concerns and clean abstractions. The stack organizes power management into a fast, low-level reactive layer that is vertically coupled to a strategic, high-level deliberative layer. An agent monitors local job performance and hardware utilization, optimizing its power requests to achieve its performance goals while accounting for global datacenter conditions and competition from other agents. The reactive policy could adjust a processor’s power mode in response to program phases while the deliberative policy ensures each processor’s adjustment anticipates other processors’ policies and the datacenter’s broader goals in sustainability, safety, and stability.
@@
The cognitive stack could leverage multi-agent game theory and reinforcement learning for dynamic decision making (yeh24, ). Dynamism is important because computation exhibits time-varying behavior, and allocation decisions in the present should account for the past and anticipate the future. For example, consider a repeated game in which agents spend tokens for power. Each agent learns a policy for spending tokens, requesting power, and mapping power to datacenter resources. When carbon-free energy is scarce, the datacenter could offer more tokens to users that defer jobs or require more tokens from those who compute. How should agents spend tokens to maximize long-term performance when allocations in one time period affect those in an uncertain future? How should the datacenter price power to achieve sustainability or DR goals?
@@
Risks. Some might wonder whether DR will be necessary given how datacenter operators are investing in renewable energy. Today, net zero claims rely on wind/solar energy investments that offset datacenter energy use, but carbon-free energy is too often generated at times and places that do not align with when and where computation happens such that ”net-zero” datacenters consume carbon-intensive energy in many hours of the year (acun23, ). Others might wonder whether DR already exists. Today’s narrow solutions focus on stability as grids request reduced use to avoid rare power emergencies, but even these suffer from incentive problems (wierman14, ). Researchers have studied how datacenters could participate in DR using simplifying assumptions. For example, studies assume 20% of power is consumed by batch jobs that are deferrable within a 24-hour window without loss of utility (wierman14, ; radovanovic21, ), but practical DR must accommodate diverse mixes of heterogeneous jobs.
@@ 5. Energy Economics
Economics and policy shape the solution space for carbon-efficient computing. Governments might implement carbon trading or incentives for low-carbon energy. The private sector might implement offset programs, leading to renewable energy purchase agreements and credits. Future demand response frameworks will require sophisticated markets that price electricity at its true social marginal cost and incentivize users to schedule computation accordingly. Although there is extensive literature on low-carbon policies for other industries (abrell19, ), economic analysis of policies specifically aimed at computing is relatively unexplored. In the near term, industry will benefit from policy-induced incentives when investing in renewable energy supply (e.g., renewable energy certificates). But as supply grows, industry must be able to monetize its flexibility in energy demand. Datacenters are often the largest consumers on the grid and we must understand how their locally optimized decisions for net zero operations will affect other consumers and impact society.
@@
Furthermore, we will study how improved sustainability impact demand for computing. Given an unpriced environmental externality (hardin68, ), such as carbon, one might ask whether society is computing too much. What is the optimal amount of computing for society? Will more efficient algorithms and systems lead to so much demand for new computing applications that overall carbon will increase? The Jevons Paradox states increased efficiency may not reduce demand for energy in the long run (and may even increase it). Prior research suggests, as a technology becomes more efficient, its use increases and produces rebound effects that range from 10% to 40%, reducing but not eliminating energy savings (gillingham16, ). But there has been no study of these effects for computing.
@@
We would need to estimate three types of rebound effects as technological efficiency lowers operating costs. First, direct effects arise when lower costs increase use of the technology. Datacenters likely exhibit strong direct effects as operators provision hardware to maximally use provisioned power (fan07, ); more efficient processors lead to datacenters with more processors. Second, indirect effects arise when lower costs increase use of other technologies. Quantifying indirect effects requires understanding substitutability and complementary between hardware components, which in turn depend on hardware capacity translates into software performance; more efficient processors may lead to datacenters with more memory as well. Finally, macroeconomic effects arise when lower costs encourage technology use for new applications. Efficient processors may scale the use of large AI models for everyday tasks (e.g., conversational bots) rather than niche tasks (e.g., playing games).